OpenStudy (anonymous):

MEDAL AND FAN Complete the equation of the graphed linear function. Write the slope in decimal form. y = x + The graph of the function goes through point negative 4, negative 1 and point 0, negative 4.

OpenStudy (anonymous):

Do you know the slope formula?

OpenStudy (anonymous):

no but i am not good at this i forgot it

OpenStudy (anonymous):

please help me

OpenStudy (anonymous):

so answer is?

OpenStudy (anonymous):

so if you plug in the x and y intercepts then you will get \[\frac{ (-4)-1 }{ 0-4 }\]

OpenStudy (anonymous):

now I do hope you can do simple subtraction because I'm not just gonna give you the answer and i'm sorry the proper equation is \[\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (anonymous):

ok i can can you set up another

OpenStudy (anonymous):

just subtract 1 from -4 giving you -5 then subtract 4 from 0 and maintain the fraction and there's your answer

OpenStudy (anonymous):

Complete the equation of the graphed linear function in point-slope form. y – (–2) = (x – ) The graph of the function goes through point 1, negative 2 and point 2, 2. http://static.k12.com/calms_media/media/1522000_1522500/1522154/1/b46c3a188a94fdd3b92220d260520511848f250c/MS_IMC-140829-100656.jpg

OpenStudy (anonymous):

the ? is - it is broke lol

OpenStudy (anonymous):

nah I can see it just refresh your page

OpenStudy (anonymous):

oh ok thanks

OpenStudy (anonymous):

\[\frac{ 2-(-2) }{ 2-1}\]

OpenStudy (anonymous):

so first one is -4 and second is 1?

OpenStudy (anonymous):

nvm positive 4

OpenStudy (anonymous):

so first 4 and second 1?

OpenStudy (anonymous):

Yessir

OpenStudy (anonymous):

THANK YOU!! I WILL CHECK SCORE

OpenStudy (anonymous):

NOOOOOOOOOOOOOOOOOOOOOOOOO I GOT 20%

OpenStudy (anonymous):

i need 80%+

OpenStudy (anonymous):

lemme guess you're using edgenuity?

OpenStudy (anonymous):

k12.com

OpenStudy (anonymous):

i am in mid school

OpenStudy (anonymous):

Pre Algebra?

OpenStudy (anonymous):

or Algebra 1?

OpenStudy (anonymous):

idk probly algebra 1 i did pre in 6th grade

OpenStudy (anonymous):

well sorry i just wiat on my mom when she gets home g2g to vocab thank you

OpenStudy (anonymous):

Anytime